OpenStudy (anonymous):

Which of the following shows 2 + (x + 3y) rewritten using the Associative Property of Addition? 2 + x + 3y (2 + x) + 3y 2x + (3 + y) x + (2 + 3y)

OpenStudy (welshfella):

Associative property:- a + (b + c) = (a + b) + c

OpenStudy (anonymous):

wat the answer

OpenStudy (welshfella):

well check the choices against my last post to see which one is similar

OpenStudy (anonymous):

none are similar

OpenStudy (welshfella):

let a = 2 , b = x and c = 3y

OpenStudy (welshfella):

can you see that the following are similar a + ( b + c) = (a + b) + c 2 + (x + 3y) =

OpenStudy (welshfella):

now what goes after the '='

OpenStudy (welshfella):

as I said a = 2 etc

OpenStudy (anonymous):